OpenStudy (anonymous):

27/28 x 7/18 = ? havinq trouble with this prob.

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

189/504 reduced = 3/8

OpenStudy (anonymous):

you just multiply them together then reduce

OpenStudy (anonymous):

(27*7)/(28*18)

OpenStudy (anonymous):

\[3^{3}*7/(2^{2}*7*2*3^{2})\] just cancel the numbers and arrive at 3/(2^2*2)=3/8